OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 5 meters and legs that measure 3 meters and 4 meters. What is the surface area of this prism if the height of the prism is 6 meters?

OpenStudy (anonymous):

area of each base = 3*4 * 1/2 = 6 sq m area of 3 sides = 5* 6 = 30 sq m, + 4*6 = 24 and 3* 6 = 18 total surface area = 2 * 6 + 30 + 24 + 18 = 12 + 30 + 24 + 18 = 84 sq m